What volume will 200 ml of gas at 20 o C and 800 mm Hg occupy at STP?


1
Expert's answer
2022-06-15T05:32:04-0400

Given:

P1 = 800 mmHg

V1 = 200 mL

T1 = 20°C + 273.15 = 293.15 K

P2 = 760 mmHg (STP)

V2 = unknown (STP)

T2 = 0°C + 273.15 = 273.15 K (STP)


Formula: P1V1/T1 = P2V2/T2


Solution:

Combined gas law can be used.

Combined gas law can be expressed as:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (800 mmHg × 200 mL × 273.15 K) / (760 mmHg × 293.15 K) = 196.2 mL

V2 = 196.2 mL


Answer: 196.2 mL